11. Potassium-40 decays to argon-40 with a half-life of 1.27x10 y. What is the age of a rock in which the mass ratio of 4°Ar to oK is 4.2? [15 pt)
Answer #1

t1/2    = 1.27*10^9 y

K     = 0.693/t/12

       = 0.693/1.27*10^9   = 5.45*10^-10 y^-1

The ration of Ar-40 to K-40 is 4.2 to 1

Then the fraction of K-40 remains   = 1/4.2+1 = 1/5.2   = 0.1923

ln[A]/[A0]     = -Kt

ln0.1923    = -5.45*10^-10*t

-1.6489     = 5.45*10^-10 * t

t          = 1.6489/5.45*10^-10    = 3.02*10^9 y >>>answer